OpenStudy (anonymous):

If f(x) = 2x + 8, find f(9).

OpenStudy (sayan):

f(9)=2*9+8 =26

OpenStudy (anonymous):

Plug in 9 for x. 2(9)+8=26=y

OpenStudy (anonymous):

thank you ((: